# Discount Model Search for Quality Diversity Optimization in High-Dimensional Measure Spaces

### v1

@@ Abstract.
Quality diversity (QD) optimization searches for a collection of solutions that optimize an objective while attaining diverse outputs of a user-specified, vector-valued measure function. Contemporary QD algorithms focus on low-dimensional measures because high-dimensional measures are prone to distortion, where many solutions found by the QD algorithm map to similar measures. For example, the CMA-MAE algorithm guides measure space exploration with a histogram in measure space that records so-called discount values. However, CMA-MAE stagnates in domains with high-dimensional measure spaces because solutions with similar measures fall into the same histogram cell and thus receive identical discount values. To address these limitations, we propose Discount Model Search (DMS), which guides exploration with a model that provides a smooth, continuous representation of discount values. In high-dimensional measure spaces, this model enables DMS to distinguish between solutions with similar measures and thus continue exploration. We show that DMS facilitates new QD applications by introducing two domains where the measure space is the high-dimensional space of images, which enables users to specify their desired measures by providing a dataset of images rather than hand-designing the measure function. Results in these domains and on high-dimensional benchmarks show that DMS outperforms CMA-MAE and other black-box QD algorithms.
quality diversity optimization, black-box optimization, derivative-free optimization, evolutionary algorithms, MAP-Elites
conference: ; ; $c$s: Computing methodologies Machine learning
1. Introduction
@@
We present a method that enhances exploration in quality diversity (QD) optimization and show how this method enables new applications for QD. QD (Pugh et al., 2016) is a branch of stochastic optimization that seeks to find diverse, high-performing solutions to a problem, with applications like robotics (Mouret and Clune, 2015), generative modeling (Ding et al., 2024), and LLM red-teaming (Samvelyan et al., 2025). To illustrate, consider searching for “a photo of a hiker.” By itself, this problem is ambiguous since hikers vary widely in appearance, depending on, for instance, where they are hiking and the time of year. If we run a single-objective optimization algorithm like Adam (Kingma and Ba, 2015) or CMA-ES (Hansen, 2016), the image we find would optimize the objective $f$ of “a photo of a hiker,” but the hiker could take on one of many different appearances. In contrast, QD (Pugh et al., 2016) can manage the ambiguity by searching for an archive (set) of images that both optimize the objective $f$ and diversify along the outputs of a measure function $\bm{m}$}.
@@
Prior work (Pugh et al., 2016; Cully et al., 2015; Fontaine and Nikolaidis, 2021) typically hand-designs $\bm{m}$} to output low-dimensional (<10D) vectors. To illustrate, our measure function could output two measures: the hiker’s age and the temperature for which they are dressed. Then, our archive would contain images like a younger hiker dressed for cold weather and an older hiker dressed for warm weather, as well as all images in between.
@@
Figure 1. (a): One failure mode of CMA-MAE. On a flat objective $f$, solutions ${\bm{\theta}}_{1}$ and ${\bm{\theta}}_{2}$ fall in the same archive cell based on their measures, resulting in identical discount values from the discount function $f_{A}$. (b): In our proposed DMS, the discount model provides a smooth discount function that assigns distinct discount values to ${\bm{\theta}}_{1}$ and ${\bm{\theta}}_{2}$, showing that ${\bm{\theta}}_{2}$ has greater archive improvement than ${\bm{\theta}}_{1}$ ($\Delta_{2}>\Delta_{1}$) and thus providing a stronger signal to guide search. (c): Number of unique cells where solutions sampled by CMA-MAE land in two benchmarks (mean over 20 trials; Sec. 4).
@@
One reason prior works focus on low-dimensional measures is that high-dimensional measure spaces are prone to distortion, where many solutions (images) map to a small region of measure space (i.e., the solutions have similar measures). For example, it may be easy to search for images of hikers in warm weather, while hikers in cold weather are hard to find. Although distortion exists in low-dimensional measure spaces  (Fontaine and Nikolaidis, 2021; Fontaine et al., 2020), it is more prominent in high-dimensional measure spaces because there are exponentially larger volumes to which solutions can map (Sec. 4).
@@
We propose to scale to high-dimensional measure spaces by addressing the effects of distortion in Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) (Fontaine and Nikolaidis, 2023), a state-of-the-art black-box QD algorithm. To fill the archive of solutions, CMA-MAE searches for solutions $\bm{\theta}$} that maximize archive improvement, defined as $\Delta({\bm{\theta}})=f({\bm{\theta}})-f_{A}({\bm{m}}(\theta))$, where $f({\bm{\theta}})$ is the solution’s objective value and $f_{A}$ is a discount function that returns scalar discount values based on the solution’s measures ${\bm{m}}({\bm{\theta}})$. CMA-MAE represents $f_{A}$ as a histogram by tessellating the measure space into cells and storing a scalar value in each cell. In domains with high distortion, particularly domains with high-dimensional measures, solutions with similar measures fall into the same cell. As a result, CMA-MAE incorrectly assigns these solutions the same discount value, which creates inaccurate improvement values that cause the search to stagnate (1(a)).
@@
Our key insight is that a smooth, continuous representation of the discount function will enhance exploration in high-dimensional measure spaces. As such, we propose Discount Model Search (DMS), where a discount model assigns distinct discount values to solutions even when distortion causes them to have similar measures (1(b)). The discount values guide DMS to explore the measure space and discover solutions long after CMA-MAE would stagnate.
@@
Figure 2. In the LSI (Hiker) domain, the objective is “A photo of the face of a hiker,” and the measure space is the space of images. We specify desired measures with landscape images from LHQ (Skorokhodov et al., 2021). Thus, DMS finds images depicting what a hiker might look like in each landscape: hikers in thick jackets for the mountains or lighter clothing for the beach, and even a baby bundled up for the snow. Each hiker is shown to the left of their corresponding landscape.
2. Problem Formulation
As formulated in Fontaine and Nikolaidis (2021), black-box quality diversity (QD) optimization considers a scalar-valued objective function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and a vector-valued measure function ${\bm{m}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$. Both functions take as input a solution ${\bm{\theta}}\in\mathbb{R}^{n}$, and $\bm{m}$} outputs $k$ measures. The image of $\bm{m}$} forms the measure space $S$. The QD objective is to find, for every ${\bm{s}}\in S$, a solution $\bm{\theta}$} such that ${\bm{m}}({\bm{\theta}})={\bm{s}}$ and $f({\bm{\theta}})$ is maximized. As stated, this QD objective requires infinite memory since $S$ is a continuous space. Hence, algorithms based on MAP-Elites (Mouret and Clune, 2015) discretize $S$ into a tessellation $T$ of $M$ cells, leading to a relaxed QD objective $\max_{{\bm{\theta}}_{1..M}}\sum_{i=1}^{M}f({\bm{\theta}}_{i})$. Each solution ${\bm{\theta}}_{i}$ has measures located in the region of measure space indicated by cell $i$ in $T$, and the set of solutions ${\bm{\theta}}_{1..M}$ is referred to as an archive $\mathcal{A}$.
3. Background
Our work builds on several black-box QD algorithms. We include further related work in QD and image generation in Sec. 8.
MAP-Elites (Mouret and Clune, 2015) produces a “grid archive” where the tessellation $T$ divides the measure space into a grid of axis-aligned (hyper-)rectangles. Each cell in the archive stores one solution. Each iteration, MAP-Elites selects an archive solution $\bm{\theta}$}, mutates it to create a new solution $\bm{\theta}$$^{\prime}$, and adds $\bm{\theta}$$^{\prime}$ to the archive. As $\bm{\theta}$$^{\prime}$ is added, it is assigned to a cell $e$ based on its measure values. $\bm{\theta}$$^{\prime}$ replaces the solution in cell $e$ if it has a higher objective value. In this manner, MAP-Elites retains elites, i.e., the best solution found in each cell. We consider a version of MAP-Elites that mutates solutions by adding isotropic Gaussian noise, i.e., new solutions are created as ${\bm{\theta}}^{\prime}\leftarrow{\bm{\theta}}+\sigma\mathcal{N}(\bm{0},\bm{I})$.
@@ 3. Background
MAP-Elites (line) (Vassiliades and Mouret, 2018) augments MAP-Elites with the Iso+LineDD operator, which leverages correlations between solutions in the archive by generating new solutions as ${\bm{\theta}}^{\prime}\leftarrow{\bm{\theta}}_{1}+\sigma_{1}\mathcal{N}(\bm{0},\bm{I})+\sigma_{2}\mathcal{N}(\bm{0},\bm{I})({\bm{\theta}}_{2}-{\bm{\theta}}_{1})$. ${\bm{\theta}}_{1}$ and ${\bm{\theta}}_{2}$ are sampled from the archive.
Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) (Fontaine and Nikolaidis, 2023) is a state-of-the-art black-box QD algorithm that integrates the CMA-ES (Hansen, 2016) optimizer into MAP-Elites to directly optimize the QD objective (Sec. 2). CMA-MAE maintains a CMA-ES instance that samples solutions ${\bm{\theta}}_{i}$ from a Gaussian distribution $\mathcal{N}({\bm{\theta}}^{*},{\bm{\Sigma}})$. Each solution ${\bm{\theta}}_{i}$ is evaluated and added to the archive. For each ${\bm{\theta}}_{i}$, CMA-MAE computes an improvement value $\Delta_{i}$ that represents how much ${\bm{\theta}}_{i}$ improves the archive. The ranking of the improvement values enables the CMA-ES instance to update the distribution parameters $\bm{\theta}$$^{*}$ and $\bm{\Sigma}$} in the direction of greatest archive improvement; thus, CMA-ES continues to sample solutions that improve the archive in future iterations. This update causes CMA-ES to optimize the QD objective. Multiple CMA-ES instances may operate in parallel, with each instance referred to as an emitter.
@@
4. Understanding Distortion in High-Dimensional Measure Spaces
Distortion in QD refers to when large areas of solution space map to a small region of measure space (Fontaine and Nikolaidis, 2023). When CMA-MAE encounters distortion, the solutions it samples land in fewer archive cells since they have similar measures. 1(a) shows one scenario where solutions that land in the same cell interfere with CMA-MAE’s improvement mechanism. On a flat (constant) objective like the one in the figure, CMA-MAE’s histogram represents how often it has visited each area of measure space, and the improvement ranking guides it towards areas that it has not visited before. For instance, Cell 3 has the lowest discount value among the three cells since it has not been explored yet, so the direction of greatest archive improvement is to generate solutions that land in Cell 3. However, since ${\bm{\theta}}_{1}$ and ${\bm{\theta}}_{2}$ both land in Cell 2, they have the same discount value. Since they also have the same constant objective, they receive the same improvement value ($\Delta_{1}=\Delta_{2}$). Hence, CMA-MAE cannot identify the direction of greatest archive improvement.
In 1(c), we present an experiment that illustrates how distortion causes solutions to land in the same cell, which activates failures like the one above. We run CMA-MAE in the 2D and 10D LP (Sphere) benchmarks, which are designed to exhibit distortion (Sec. 6.1). 2D and 10D indicate the measure space dimensionality. We plot the number of unique archive cells where solutions sampled by CMA-MAE land according to their measures; CMA-MAE samples 540 solutions per iteration. The plot shows that in both benchmarks, CMA-MAE begins searching in areas of low distortion, as many solutions land in unique cells. Over time, solutions often land in the same cell (the number of unique cells goes down), indicating CMA-MAE has reached areas with higher distortion.
@@ 4. Understanding Distortion in High-Dimensional Measure Spaces
Since large cells can amplify the effects of distortion, Fontaine and Nikolaidis (2023) suggest increasing the archive resolution (i.e., adding more cells), albeit only in a 2D measure space. With higher resolution, cells are smaller, so solutions with similar measures can still fall in different cells and receive different discount values. However, this approach entails large amounts of memory, and this amount grows exponentially with measure space dimensionality. Since increasing the archive resolution effectively makes the histogram closer to a continuous function, we propose to eliminate the histogram entirely and instead search with a continuous representation of the discount function.
5. Discount Model Search
@@ 44      Regress $\hat{f}_{A}$ on $\mathcal{D}_{A}$
(1) $t_{A}=\begin{cases}\hat{f}_{A}({\bm{s}})&\text{if }f({\bm{\theta}})\leq\hat{f}_{A}({\bm{s}})\\ (1-\alpha)\hat{f}_{A}({\bm{s}})+\alpha f({\bm{\theta}})&\text{if }f({\bm{\theta}})>\hat{f}_{A}({\bm{s}})\end{cases}$
@@
6. Domains
We evaluate DMS on standard QD benchmarks and in a setting we refer to as Quality Diversity with Datasets of Measures (QDDM). We provide further details of all domains in Appendix D.
6.1. Benchmarks

### v2

@@ Abstract
Quality diversity (QD) optimization searches for a collection of solutions that optimize an objective while attaining diverse outputs of a user-specified, vector-valued measure function. Contemporary QD algorithms are typically limited to low-dimensional measures because high-dimensional measures are prone to distortion, where many solutions found by the QD algorithm map to similar measures. For example, the state-of-the-art CMA-MAE algorithm guides measure space exploration with a histogram in measure space that records so-called discount values. However, CMA-MAE stagnates in domains with high-dimensional measure spaces because solutions with similar measures fall into the same histogram cell and hence receive the same discount value. To address these limitations, we propose Discount Model Search (DMS), which guides exploration with a model that provides a smooth, continuous representation of discount values. In high-dimensional measure spaces, this model enables DMS to distinguish between solutions with similar measures and thus continue exploration. We show that DMS facilitates new capabilities for QD algorithms by introducing two new domains where the measure space is the high-dimensional space of images, which enables users to specify their desired measures by providing a dataset of images rather than hand-designing the measure function. Results in these domains and on high-dimensional benchmarks show that DMS outperforms CMA-MAE and other existing black-box QD algorithms.
1 Introduction
@@
We present a method that enhances exploration in quality diversity (QD) optimization and show how this method enables new applications for QD. QD (Pugh et al., 2016) is a branch of stochastic optimization that seeks to find diverse, high-performing solutions to a problem, with applications like robotics (Mouret & Clune, 2015), generative modeling (Ding et al., 2024), and LLM red-teaming (Samvelyan et al., 2025). To illustrate, consider searching for “a photo of a hiker.” By itself, this problem is ambiguous since hikers vary widely in appearance, depending on, for instance, where they are hiking and the time of year. If we run a single-objective optimization algorithm like Adam (Kingma & Ba, 2015) or CMA-ES (Hansen, 2016), the image we find would optimize the objective $f$ of “a photo of a hiker,” but the hiker could take on one of many different appearances. In contrast, QD (Pugh et al., 2016) can manage the ambiguity by searching for an archive (set) of images that both optimize the objective $f$ and diversify along the outputs of a measure function $\bm{m}$}.
@@
Prior work (Pugh et al., 2016) typically hand-designs $\bm{m}$} to output low-dimensional (<10D) vectors. To illustrate, our measure function could output two measures: the hiker’s age and the temperature for which they are dressed. Then, our archive would contain images like a younger hiker dressed for cold weather and an older hiker dressed for warm weather, as well as all images in between.
@@
Figure 1: (a): One failure mode of CMA-MAE. On a flat objective $f$, solutions ${\bm{\theta}}_{1}$ and ${\bm{\theta}}_{2}$ fall in the same archive cell based on their measures, resulting in identical discount values from the discount function $f_{A}$. (b): In our proposed DMS, the discount model provides a smooth discount function that assigns distinct discount values to ${\bm{\theta}}_{1}$ and ${\bm{\theta}}_{2}$, showing that ${\bm{\theta}}_{2}$ has greater archive improvement than ${\bm{\theta}}_{1}$ ($\Delta_{2}>\Delta_{1}$) and thus providing a stronger signal to guide search. (c): Number of unique cells where solutions sampled by CMA-MAE land in two benchmarks (mean over 20 trials; Sec. 4).
@@
One reason prior works focus on low-dimensional measures is that high-dimensional measure spaces are prone to distortion, where many solutions (images) map to a small region of measure space (i.e., the solutions have similar measures). For example, it may be easy to search for images of hikers in warm weather, while hikers in cold weather are hard to find. Although distortion exists in low-dimensional measure spaces  (Fontaine & Nikolaidis, 2021; Fontaine et al., 2020), it is more prominent in high-dimensional measure spaces because there are exponentially larger volumes to which solutions can map (Sec. 4).
@@
We propose to scale to high-dimensional measure spaces by addressing the effects of distortion in Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) (Fontaine & Nikolaidis, 2023), a state-of-the-art black-box QD algorithm. To fill the archive of solutions, CMA-MAE searches for solutions $\bm{\theta}$} that maximize archive improvement, defined as $\Delta({\bm{\theta}})=f({\bm{\theta}})-f_{A}({\bm{m}}(\theta))$, where $f({\bm{\theta}})$ is the solution’s objective value and $f_{A}$ is a discount function that returns scalar discount values based on the solution’s measures ${\bm{m}}({\bm{\theta}})$. CMA-MAE represents $f_{A}$ as a histogram by tessellating the measure space into cells and storing a scalar value in each cell. In domains [^1] In our work, the term “domain” is synonymous with “problem.” with high distortion, particularly domains with high-dimensional measures, solutions with similar measures fall into the same cell. As a result, CMA-MAE incorrectly assigns these solutions the same discount value, which creates inaccurate improvement values that cause the search to stagnate (1(a)).
@@
Our key insight is that a smooth, continuous representation of the discount function will enhance exploration in high-dimensional measure spaces. As such, we propose Discount Model Search (DMS), where a smooth discount model accurately assigns distinct discount values to solutions even when distortion causes them to have similar measures (1(b)). The discount values guide DMS to explore the measure space and discover solutions long after CMA-MAE would stagnate.
@@
Figure 2: In the LSI (Hiker) domain, the objective is “A photo of the face of a hiker,” and the measure space is the space of images. We specify desired measures with landscape images from LHQ (Skorokhodov et al., 2021). Thus, DMS finds images depicting what a hiker might look like in each landscape: hikers in thick jackets for the mountains or lighter clothing for the beach, and even a baby bundled up for the snow. Each hiker is shown to the left of their corresponding landscape.
2 Problem Formulation
As formulated in Fontaine & Nikolaidis (2021), black-box quality diversity (QD) optimization considers a scalar-valued objective function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and a vector-valued measure function ${\bm{m}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$. Both functions take as input a solution ${\bm{\theta}}\in\mathbb{R}^{n}$, and $\bm{m}$} outputs $k$ measures. The image of $\bm{m}$} forms the measure space $S$. The QD objective is to find, for every ${\bm{s}}\in S$, a solution $\bm{\theta}$} such that ${\bm{m}}({\bm{\theta}})={\bm{s}}$ and $f({\bm{\theta}})$ is maximized. As stated, this QD objective requires infinite memory since $S$ is a continuous space. Hence, algorithms based on MAP-Elites (Mouret & Clune, 2015) discretize $S$ into a tessellation $T$ of $M$ cells, leading to a relaxed QD objective $\max_{{\bm{\theta}}_{1..M}}\sum_{i=1}^{M}f({\bm{\theta}}_{i})$. Each solution ${\bm{\theta}}_{i}$ has measures located in the region of measure space indicated by cell $i$ in $T$, and the set of solutions ${\bm{\theta}}_{1..M}$ is referred to as an archive $\mathcal{A}$.
3 Background
Our work builds on several black-box QD algorithms. We include further related work in QD and image generation in Appendix A.
MAP-Elites (Mouret & Clune, 2015) produces a “grid archive” where the tessellation $T$ divides the measure space into a grid of axis-aligned (hyper-)rectangles. Each cell in the archive stores one solution. Each iteration, MAP-Elites selects an archive solution $\bm{\theta}$}, mutates it to create a new solution $\bm{\theta}$$^{\prime}$, and adds $\bm{\theta}$$^{\prime}$ to the archive. As $\bm{\theta}$$^{\prime}$ is added, it is assigned to a cell $e$ based on its measure values. $\bm{\theta}$$^{\prime}$ replaces the solution in cell $e$ if it has a higher objective value. In this manner, MAP-Elites retains elites, i.e., the best solution found in each cell. We consider a version of MAP-Elites that mutates solutions by adding isotropic Gaussian noise, i.e., new solutions are created as ${\bm{\theta}}^{\prime}\leftarrow{\bm{\theta}}+\sigma\mathcal{N}(\bm{0},\bm{I})$.
@@ 3 Background
MAP-Elites (line) (Vassiliades & Mouret, 2018) augments MAP-Elites with the Iso+LineDD operator, which leverages correlations between solutions in the archive by generating new solutions as ${\bm{\theta}}^{\prime}\leftarrow{\bm{\theta}}_{1}+\sigma_{1}\mathcal{N}(\bm{0},\bm{I})+\sigma_{2}\mathcal{N}(\bm{0},\bm{I})({\bm{\theta}}_{2}-{\bm{\theta}}_{1})$. ${\bm{\theta}}_{1}$ and ${\bm{\theta}}_{2}$ are sampled from the archive.
Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) (Fontaine & Nikolaidis, 2023) is a state-of-the-art black-box QD algorithm that integrates the CMA-ES (Hansen, 2016) optimizer into MAP-Elites to directly optimize the QD objective (Sec. 2). CMA-MAE maintains a CMA-ES instance that samples solutions ${\bm{\theta}}_{i}$ from a Gaussian distribution $\mathcal{N}({\bm{\theta}}^{*},{\bm{\Sigma}})$. Each solution ${\bm{\theta}}_{i}$ is evaluated and added to the archive. For each ${\bm{\theta}}_{i}$, CMA-MAE computes an improvement value $\Delta_{i}$ that represents how much ${\bm{\theta}}_{i}$ improves the archive. The ranking of the improvement values enables the CMA-ES instance to update the distribution parameters $\bm{\theta}$$^{*}$ and $\bm{\Sigma}$} in the direction of greatest archive improvement; thus, CMA-ES continues to sample solutions that improve the archive in future iterations. This update causes CMA-ES to optimize the QD objective. Multiple CMA-ES instances may operate in parallel, with each instance referred to as an emitter.
@@
4 Understanding Distortion in High-Dimensional Measure Spaces
Distortion in QD refers to when large areas of solution space map to a small region of measure space (Fontaine & Nikolaidis, 2023). When CMA-MAE encounters distortion, the solutions it samples land in fewer archive cells since they have similar measures. 1(a) shows one scenario where solutions that land in the same cell interfere with CMA-MAE’s improvement mechanism. On a flat (constant) objective like the one in the figure, CMA-MAE’s histogram represents how often it has visited each area of measure space, and the improvement ranking guides it towards areas that it has not visited before. For instance, Cell 3 has the lowest discount value among the three cells since it has not been explored yet, so the direction of greatest archive improvement is to generate solutions that land in Cell 3. However, since ${\bm{\theta}}_{1}$ and ${\bm{\theta}}_{2}$ both land in Cell 2, they have the same discount value. Since they also have the same constant objective, they receive the same improvement value ($\Delta_{1}=\Delta_{2}$). Hence, CMA-MAE cannot identify the direction of greatest archive improvement.
In 1(c), we present an experiment that illustrates how distortion causes solutions to land in the same cell, which activates failures like the one above. We run CMA-MAE in the 2D and 10D LP (Sphere) benchmarks, which are designed to exhibit distortion (Sec. 6.1). 2D and 10D indicate the measure space dimensionality. We plot the number of unique archive cells where solutions sampled by CMA-MAE land according to their measures; CMA-MAE samples 540 solutions per iteration. The plot shows that in both benchmarks, CMA-MAE begins searching in areas of low distortion, as many solutions land in unique cells. Over time, solutions more often land in the same cell (i.e., the number of unique cells goes down), indicating CMA-MAE has reached areas with higher distortion.
@@ 4 Understanding Distortion in High-Dimensional Measure Spaces
Since large cells can amplify the effects of distortion, Fontaine & Nikolaidis (2023) suggest increasing the archive resolution (i.e., adding more cells), albeit only in a 2D measure space. With higher resolution, cells are smaller, so solutions with similar measures can still fall in different cells and receive different discount values. However, this approach entails large amounts of memory, and this amount grows exponentially with measure space dimensionality. Since increasing the archive resolution effectively makes the histogram closer to a continuous function, we propose to eliminate the histogram entirely and instead search with a continuous representation of the discount function.
5 Discount Model Search
@@ 44      Regress $\hat{f}_{A}$ on $\mathcal{D}_{A}$
$t_{A}=\begin{cases}\hat{f}_{A}({\bm{s}})&\text{if }f({\bm{\theta}})\leq\hat{f}_{A}({\bm{s}})\\ (1-\alpha)\hat{f}_{A}({\bm{s}})+\alpha f({\bm{\theta}})&\text{if }f({\bm{\theta}})>\hat{f}_{A}({\bm{s}})\end{cases}$ (1)
@@
6 Domains
We evaluate DMS on standard QD benchmarks and in a setting we refer to as Quality Diversity with Datasets of Measures (QDDM). We provide further details of all domains in Appendix E.
6.1 Benchmarks